# Repetition Improves Language Model Embeddings

### v1

@@ 3.3 Repetition Captures Bidirectional Info
With classical embeddings, the contextualized embeddings of $A$ parts of $s^{+},s^{-},q$ are identical by construction. To test whether echo embeddings can meaningfully distinguish $s^{+}$ and $s^{-}$ despite having identical initial tokens, we take the mean over just the $A$-portion of the echo embeddings and plot the cosine similarities $\operatorname{Sim}(q,s^{+})$ and $\operatorname{Sim}(q,s^{-})$ in Figure 2 (left). We find that $\operatorname{Sim}(q,s^{+})$ is typically larger than $\operatorname{Sim}(q,s^{-})$. Since we are only pooling the echo embeddings of the A-portion, any distinction between $s^{+},s^{-}$ must come from the echo embeddings of $A$ capturing information from the later parts of the sentence. This showcases that current autoregressive language models can in fact allow early tokens to capture information from later tokens via echo embeddings.
3.4 Classical vs. Echo on Synthetic Data
In Section 3.3 we demonstrated that echo embeddings encode bidirectional information. However, is this sufficient to recover from the failure mode of classical embeddings? Further, where will we expect echo embeddings to improve over classical embeddings? Here, we compare echo and classical embeddings on synthetic data to answer both of these questions.
Datasets.
We sample datasets according to two structures depending on whether the discriminating information between $s^{+}$ and $s^{-}$ is in the first half (structure S1) or second half (structure S2) of the sentence. Using the structures below, we generate samples using GPT-4, as in the previous section (full details in the appendix):
(S1) $q\colon[A,B]$, $s^{+}\colon[A^{+},B^{+}]$, $s^{-}\colon[A^{+},B^{-}]$
(S2) $q\colon[A,B]$, $s^{+}\colon[A^{+},B^{+}]$, $s^{-}\colon[A^{-},B^{+}]$.
We measure the accuracy of identifying which of two sentences $s^{+}$ and $s^{-}$ is closer to the query as measured by the cosine similarity in the embeddings. We compare classical vs echo embeddings when using mean pooling to aggregate embeddings extracted from Mistral-7B model. We use mean token embedding rather than last token because last token embeddings can be quite fragile in a zero-shot setting (Section 5.1).
Results.
We present results on the two different structures in Figure 2 (right). We see that classical embeddings struggle on Structure 1—when the distinguishing information is at the beginning, the embeddings corresponding to these early tokens exaggerate similarity between $q$ and $s^{-}$ because they do not encode the information provided by $B^{-}$. In contrast, echo embeddings are able to successfully determine the more similar sentence, presumably because embeddings of the $A$-portion now also encode information about the later parts (demonstrated in Section 3.3). As a control for other reasons echo embeddings outperform classical embeddings, we also compare these embeddings on structure two, where early tokens provide discriminative signal without needing the later context. As expected, both classical and echo embeddings achieve good performance in this setting. All in all, this analysis on synthetic data demonstrates that zero-shot classical embeddings do not encode information about later context in early token embeddings, but echo embeddings can do so.
Does last-token pooling resolve the failure of classical embeddings?
The embedding of the last token $\phi_{-1}(x)$ is, in principle, can encode information from the entire input. However, we posit that the last-token pooling strategy is highly brittle and can depend too strongly on the tokens near the end of the input. To verify this, we compare the accuracy of mean token pooling and last-token pooling for classical and echo embeddings in two settings. First, we evaluate on the original synthetic data of Structure 1 (Figure 3, left). Second, we evaluate on the same data, but where we append a uniformly randomly selected token to the end of each example (Figure 3, right). While last-token pooling has high accuracy on the original toy data (though still lower accuracy than echo embeddings), it fails to perform well on the noisy examples. Echo embeddings with mean token pooling, however, are robust to the noise.
While this particular distribution of noise is artificial, it highlight that last-token pooling can be sensitive to noise in the last token. We verify in Section 5.1 that last-token embeddings perform poorly on real data. Thus, even if last-token embeddings address the inability of mean token classical embeddings to encode information from tokens that appear later in the sequence, they are not practical due to their sensitivity to noise.
Does last-token pooling resolve the failure after finetuning?
In practice, it is common to finetune embeddings on a sentence similarity objective. It is hard to delineate the degree to which this failure mode remains after finetuning. Nonetheless, we demonstrate in Section 5.2 that our method improves in the finetuning setting, even when using last-token pooling.
Figure 3: We compare the accuracy of classical and echo embeddings, and mean and last-token pooling on sentences which have similar beginnings (Structure 1). We plot these accuracies on the original data (left), and the data in which a single uniformly randomly chosen token is appended to the end of each sentence (right).
4 Methodology
In Section 3, we explored how echo embeddings can improve over classical embeddings by addressing a fundamental failure mode. In this section, we describe the methodology by which we evaluate echo embeddings on large scale real datasets in both the zero-shot and finetuning settings. While finetuning is currently necessary to achieve state-of-the-art performance, zero-shot embeddings have the advantage that they do not require expensive finetuning on top of a pretrained language model. Zero-shot results can also more clearly show how different embedding strategies work on real datasets.
4.1 Constructing Zero-shot Embeddings.
We extract zero-shot embeddings via different strategies from three language models: Mistral-7B, LLaMA-2-7B, and LLaMA-2-13B. We select the instruction-finetuned model for each of them. Refer to Appendix A.2 for additional information on the base models. Recent literature suggests that the performance of language models on zero-shot tasks can be highly variable depending on the exact wording and template of the prompts (Sclar et al., 2023). Thus, for each of the embedding strategies we consider, we perform prompt randomization where we sample prompts by randomizing the exact wording, punctuation, and capitalization of the prompt. We describe the sampling process and the exact prompts that we use in Appendix C.
Baselines.
We compare our proposed echo embeddings (Section 3.2) to classical embeddings and two additional baselines:
  • Last-token embeddings: We mentioned in the Section 3 that last-token embeddings tend to underperform in comparsion to mean token embeddings, and thus we compare on real data.
  • Summarization: We also compare zero-shot embeddings obtained via the strategy proposed by Jiang et al. (2023b). Here, they instruct the model to summarize the input in a single word and then take the last token embedding $\phi_{-1}(x)$ as the pooled embedding of the sentence.
4.2 Constructing Finetuned Embeddings.
We adopt the conventional sentence embedding training setup (Reimers and Gurevych, 2019) where we train with a contrastive learning objective to encourage the embeddings of similar text to be close. We extract embeddings in a slightly different fashion compared to the zero-shot setting above in order to keep the finetuning methodology as similar as possible to the existing literature.
Extracting embeddings.
The training and evaluation data is separated into two categories: symmetric data, in which sentences are drawn from a single distribution (such as for sentence similarity), and asymmetric, in which the data consists of both queries and documents (such as for retrieval). We adopt a separate prompt for symmetric inputs and queries, and for documents. We construct classical embeddings by encoding text $S$ using the following prompts:
Queries & Symm. Documents
Instruct: {instruction}
@@ 4 Methodology
For echo embeddings, we use the prompts, where $S$ represents the input and $S^{\prime}=S$:
Queries & Symm. Documents
Instruct: {instruction}
$\text{Query: }S$ $\text{Query again: }S^{\prime}$
$\text{Document: }S$ $\text{Document again: }S^{\prime}$
In this case, {instruction} refers to the task instruction, which specifies a description of the task that the embedding will be used for. We adopt the instructions from Wang et al. (2023), and provide a list of the instructions in Appendix D. We append an end-of-sentence token to the end of each input, and we allow the input embedding of this token to be trainable.
Datasets.
We train on a collection of publicly available datasets that encompass both symmetric and asymmetric data that are standard training datasets in the embedding literature. We list and describe each of the datasets in Appendix D.
Optimization.

### v2

@@ 3.2 Echo Embeddings: A General Framework
In particular, let us focus on structure S2 where the discriminatory information appears in the later tokens. Echo embeddings with mean-token pooling work well, while classical embeddings did not because the early tokens could not attend to the later ones. This implies that echo embeddings can successfully address this failure and the representations of early tokens can capture information “bidirectionally” about the entire input without changing causal to bidirectional attention. To test this further, we measure whether echo embeddings can meaningfully distinguish $s^{+}$ and $s^{-}$ using just the information in the $A$ portions despite the discriminatory information solely appearing in the later $B$- parts. Recall that echo embeddings extract embeddings from the second occurrence. We plot the cosine similarities $\operatorname{Sim}(q,s^{+})$ and $\operatorname{Sim}(q,s^{-})$ in Figure 2 (C). We find that $\operatorname{Sim}(q,s^{+})$ is typically larger than $\operatorname{Sim}(q,s^{-})$. Since we are only pooling the echo embeddings of the $A$ portion, any distinction between $s^{+},s^{-}$ must come from the echo embeddings of $A$ capturing information from the later parts of the sentence.
Summary. We construct a synthetic dataset that showcases a failure mode in extracting embeddings from causal language models in the classical manner. While ours is a simplistic construction, we demonstrate that the failure mode we outline does occur on real data as well in Appendix C and could explain the poor performance of classical embeddings in practice. However, echo embeddings, where we repeat the input and take embeddings from the second occurrence, can successfully address this failure mode. We now test echo embeddings on a variety of real-world settings.
4 Experiments
Table 1: Main results in the zero-shot setting: Apples-to-apples comparison of zero-shot echo and classical embeddings with mean pooling, and other baselines on MTEB (56 datasets). For the autoregressive embeddings in this table, we extract embeddings from the Mistral-7B-Instruct-v0.1 model. Bolding only zero-shot models.
Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Average
# of datasets 12 11 3 4 15 10 1 56
Echo (Ours) 72.01 33.80 72.41 47.56 20.82 73.74 30.70 48.64
Echo (compute matched) 71.63 33.51 72.31 47.43 22.85 73.64 31.02 49.02
Classical 64.87 34.14 57.71 43.60 18.26 57.07 27.02 42.38
PromptEOL 73.84 27.53 55.47 48.44 13.18 67.14 28.33 43.69
BERT 61.47 29.48 56.33 43.44 10.44 52.98 29.82 37.87
RoBERTa 62.63 29.05 56.95 41.92 8.62 55.24 28.64 37.86
LLM2Vec (requires fine tuning) 72.51 40.06 70.95 50.43 19.74 71.90 27.84 49.43
In this section, we describe how we implement and evaluate echo embeddings on real data.
4.1 Massive Text Embedding Benchmark
MTEB & MTEB-MINI. Our main evaluation dataset is the English-language subset of the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022). MTEB is a collection of 56 datasets that are grouped into different embedding tasks: classification, clustering, pair classification, reranking, retrieval, sentence similarity (STS), and summarization, with the goal of evaluating embeddings broadly. However, as the name suggests, MTEB is massive, and requires multiple days to evaluate a 7B-parameter model on 8 A100 GPUs. Thus, we adopt a smaller 28-dataset subset of MTEB that spans all categories except summarization, which we call MTEB-MINI. We provide a detailed description of MTEB and MTEB-MINI in Appendix A.1.
4.2 Evaluation of Zero-shot Embeddings
We begin with the zero-shot setting, where embeddings are extracted from a pre-trained model without additional training. Our goal is to show that without any additional training, echo embeddings outperform prior entirely-zero-shot methods, and can perform similarly to unsupervised approaches (which require additional fine-tuning with unsupervised data), such as LLM2Vec-unsupervised (BehnamGhader et al., 2024). While supervised fine-tuning is often used to improve embedders, zero-shot embeddings have the advantage that no further (potentially expensive) training is required.
Table 2: Role of scale and base model in the zero-shot setting: Average MTEB score (56 datasets) for LLaMA-7B-2-Instruct and S-LLaMA-1.3B.
Model Echo Classical PromptEOL
S-LLaMA-1.3B 40.45 34.31 37.50
LLaMA-7B 45.84 40.29 42.84
Table 3: Casting to bidirectional attention: We evaluate MTEB-MINI (28 datasets) and compare classical with causal attention (Cl.+Uni) and with bidirectional attention (Cl.+Bi).
Model Echo Cl.+Uni Cl.+Bi
LLaMA-2-7B 56.99 47.27 43.03
S-LLaMA-1.3B 49.55 40.14 35.77
Mistral-7B 59.78 49.03 58.24
Constructing zero-shot echo embeddings. Following the echo embeddings method described in Section 3, to construct echo embeddings we need to pick a prompt, a pooling strategy, and a language model. We evaluate the impact of these choices on the performance of the embeddings on the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022).
Choosing a prompt: As specified in Section 3, we must prompt the language model to act as an autoencoder. Sclar et al. (2023) argues that on many tasks, the exact formatting and wording of the prompt can strongly influence performance. Thus, we investigate: does the choice of prompt affect embeddings quality? As we shall discuss later, we find that, the choice of prompt does not substantially affect the performance of echo embeddings (Section 4.3). This implies that prompt tuning is unnecessary, and any reasonable prompt will perform well. For our main results, we adopt the prompt “Rewrite the following paragraph: $S$. The rewritten paragraph: SS” for echo embeddings and “Write a paragraph: SS” for classical embeddings.
Table 4: Role of pooling in the zero-shot setting: Average MTEB Score (56 datasets) for mean and last token pooling.
Mean Last-token
Echo 48.64 31.55
Classical 42.38 31.94
Choosing a pooling strategy: We evaluate both mean and last-token pooling for both classical and echo embeddings 4. Consistent with our toy-data analysis in Section 3, we find that mean token pooling is required for good embeddings and thus use this for all further zero-shot experiments (Table 4).
Choosing a language model: We evaluate the performance of echo embeddings on three language models: Mistral-7B-Instruct-v0.1 (Jiang et al., 2023a), LLaMA-2-7B-Instruct (Touvron et al., 2023a), and S-LLaMA-1.3B (Xia et al., 2023) (sometimes abbreviated). We select the instruction-finetuned model for each of them. Refer to Appendix A.2 for additional information on the base models.
Figure 3: Role of prompt in the zero-shot setting: Variance of zero-shot performance across different prompts and formats on MTEB-MINI (28 datasets). Refer to Appendex D for details on the exact prompts and formats.
Compute-matching echo embeddings. Naively, echo embeddings would require twice the compute of classical embeddings. To account for this, we also consider compute-matched echo embeddings in which we halve the length of the input.
Alternative approaches to constructing embeddings from autoregressive models. In addition to our results, we compare to two recently proposed methods to convert autoregressive models into embedders without supervised fine-tuning: PromptEOL (zero-shot) and the unsupervised variant of LLM2Vec (requires additional fine-tuning). PromptEOL prompts the model to summarize the input in a single token, and extracts embeddings from the predicted token (Jiang et al., 2023b) (see Appendix D for additional details). The unsupervised variant of LLM2Vec (LLM2Vec-unsupervised) incorporates an additional masked-language-modeling fine-tuning step. Importantly, LLM2Vec is not zero-shot because it involves this extra fine-tuning step.
4.3 Zero-shot Results
Echo embeddings yield strong embeddings, and without additional compute. We present the main zero-shot results in which we evaluate echo embeddings on MTEB in Table 1. We find that echo embeddings successfully converts causal language models into strong encoders. For Mistral-7B-Instruct, we find, surprisingly, that compute-matched echo embeddings, achieves slightly higher performance than full-compute echo embeddings. The difference is small enough that it is unclear to what degree we would observe this improvement in other settings, but suggests that echo embeddings can be adapted to outperform classical embeddings without any additional compute. We extend this analysis of compute in Section 4.6. We also evaluate echo embeddings over multiple models and scales in Table 3 and find consistent results.
Comparison to PromptEOL. We find that echo embeddings outperform PromptEOL on average across the tested MTEB datasets by 5%. In addition to this improvement, echo embeddings is additionally much less sensitive to the exact wording and formatting of the prompt (Figure 3).
Comparison to prior encoder models. We evaluate the zero-shot MTEB performance of existing bidirectional (encoder) models, BERT-large and RoBERTa-large, which are often used as the backbone of fine-tuned embedders. We find that echo embeddings outperform both BERT-large and RoBERTa-large in the zero-shot setting. We suspect that the performance improvement is due to Mistral-7B being significantly more powerful than BERT and RoBERTa. This highlights the importance of our method: by using echo embeddings, we can leverage the full power of autoregressive language models, which are generally dominant over bidirectional models on other NLP tasks.
Comparison to LLM2Vec. We find that entirely zero-shot echo embeddings can perform similarly to LLM2Vec-unsupervised, despite the fact that LLM2Vec-unsupervised is not a zero-shot method, and requires additional fine-tuning on top of the pre-trained base model. Echo embeddings thus offers a simple alternative to LLM2Vec that does not require additional training.
Relaxing to bidirectional attention. We find that changing the attention mask from causal to bidirectional often substantially underperforms the causal architecture, but not for Mistral-7B (Table 3). Regardless, echo embeddings consistently outperforms this baseline. A priori, we might expect that, without further fine-tuning, such a modification to the architecture would harm performance, as we would expect that the model has not been trained to accommodate the additional attention. For LLaMA-2-7B and S-LLaMA-1.3B, this is the case. However, for Mistral-7B, casting to bidirectional attention does not harm performance and instead yields performance similar to echo embeddings. We suspect that this may arise from non-standard pre-training methodology, as discussed by BehnamGhader et al. (2024).
Ablations. We perform multiple ablations to evaluate the role of each component:
  1. 1.
    We test the role of pooling strategy in Table 4. Consistent with our analysis in Section 3, we find that mean pooling outperforms last-token pooling by a wide margin for both echo and classical embeddings. This suggests that mean pooling is required for echo embeddings in the zero-shot setting.
  2. 2.
    We evaluate the role of the exact wording and formatting of the prompt in Figure 3. We find that the performance of echo embeddings has low variance with respect to the exact choice of prompt. This implies that echo embeddings does not require prompt tuning to perform well. Further, we find that all tested echo embeddings prompts outperform every tested classical embeddings prompt.
4.4 Evaluation of Fine-tuned Embeddings
Echo embeddings also perform well in the fine-tuning setting. We will show that echo embeddings can outperform classical embeddings, even when the attention is converted to be bidirectional. Further, echo embeddings can match the performance of other methods that require much more complicated fine-tuning methodology, with identical fine-tuning and inference-time compute, and without modifying the architecture.
Table 5: Main fine-tuning results: We report MTEB scores for echo embeddings with full compute and compute-matched and classical embeddings with a causal and with a casted bidirectional architecture. For echo embeddings, we use Mistral-7B as the backbone model, and last-token pooling. For the compute matched echo embeddings, FT+IT refers to both fine-tuning and inference-time compute matching.
Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Average
# of datasets 12 11 3 4 15 10 1 56
Echo (ours) 77.43 46.32 87.34 58.14 55.52 82.56 30.73 64.68
Echo (FT+IT compute matched) 77.39 46.27 87.49 58.08 55.07 83.18 30.73 64.66
Classical 76.57 45.78 86.37 56.71 54.87 82.03 31.02 63.98
Classical + bidirectional attn. 76.70 45.94 88.15 57.23 54.96 82.42 29.32 64.23
Constructing fine-tuned echo embeddings. To construct fine-tuned embeddings, we must again choose a prompt, a pooling strategy, and a language model. In addition, we need to choose a contrastive learning objective and a collection of training datasets.
Choosing a prompt: We adopt a minimal prompt that differs for queries and documents. Prior work has shown that the addition of an instruction can improve the quality of the embeddings (Wang et al., 2023), and thus we include an instruction in the prompt. The prompts are: Echo (queries) Echo (docs) Classical (queries) Classical (docs) Instruct: {instruction}
$\text{Query: }S$ $\text{Query again: }S^{\prime}$
$\text{Document: }S$ $\text{Document again: }S^{\prime}$ Instruct: {instruction}
@@ 4.4 Evaluation of Fine-tuned Embeddings
We provide a list of the instructions in Appendix F.
Choosing a pooling strategy: We evaluate both mean and last-token pooling for both classical and echo embeddings. In practice, we find that last-token pooling performs slightly better than mean pooling in this setting (see Section 4.5).
Choosing a language model: We train and evaluate the performance of echo embeddings on Mistral-7B-Instruct-v0.1 and S-LLaMA-1.3B (Jiang et al., 2023a; Xia et al., 2023). We choose Mistral-7B as the best-performing model in the zero-shot setting, and S-LLaMA-1.3B to test the broad applicability of our method across model class and scale.
Table 6: Role of base model in the fine-tuning setting: We compare echo and classical embeddings for S-LLaMA-1.3B. FT+IT refers to the setting in which echo embeddings are compute-matched at train and inference time.
Methods MTEB Score
Echo 62.01
Echo (FT+IT compute matched) 61.65
Classical 61.26
Table 7: Comparisons to recent baselines: MTEB score for models trained on similar data. Echo embeddings matches the performance of models trained on similar data, but without requiring bidirectional attention. FT+IT refers to the setting in which echo embeddings are compute-matched at train and inference time.
Method MTEB Score
Echo (FT+IT compute matched) 64.66
LLM2Vec-supervised 64.80
GritLM-7B (public data) 64.70